Scientists once postulated the existence of a planet called Vulcan to explain


Mercury’s precession. Vulcan is supposed to be between Mercury and the sun


with the solar distance equal to 2/3 that of mercury. What would be its supposed


period?

1
Expert's answer
2021-11-14T17:12:06-0500

"\frac{T_m^2}{T_v^2}=\frac{a_m^3}{a_v^3}\\"

"T_v=\sqrt {T^2_m\frac{a_v^3}{a_m^3}}=T_m\sqrt(\frac{2a_m}{3a_m})^3 \\
T_v=87.97\sqrt(\frac{2}{3})^3=58.65 \;d."